OpenStudy (anonymous):

If a transversal intersects two parallel lines, then _____ angles are supplementary. alternate interior same-side interior acute corresponding

OpenStudy (aroub):

same-side interior

OpenStudy (anonymous):

Choose the best description for this polygon. all 8 sides are equal convex quadrilateral convex octagon concave octagon concave quadrilateral

OpenStudy (aroub):

convex octagon

OpenStudy (anonymous):

Which is the equation of the line with slope 3 that contains point (-1, 5)? y – 5 = 3(x + 1) y – 1 = 3(x + 5) y – 1 = 3(x – 5) y – 5 = 3(x - 1)